# Omni-Video: Democratizing Unified Video Understanding and Generation

### v3

@@ 3.1 Overview
Recall that our aim is to build a unified architecture that enables multimodal video understanding and generation simultaneously. To achieve this, we teach multimodal understanding models (e.g.,VILA lin2024vila ) to produce visual tokens as the input of a text-to-video diffusion model with lightweight adapters and a novel three-stage training scheme. In this way, the alignment between the output visual tokens and the diffusion model’s context conditions could be achieved efficiently, without requiring training the whole system from scratch. Fig. 2 presents the overall structure of our Omni-Video. Overall, our model connects an AR-based MLLM model for understanding and a DiT-based diffusion model for visual generation, more details are presented below.
@@ 3.2 Producing Visual Tokens with MLLMs
More specifically, the text token performs standard language modeling, predicting tokens from a vocabulary augmented with four newly introduced special tokens: <BOI>, <EOI>, <BOV>, and <EOV>. Conversely, the vision head maps the MLLMs’ hidden states into a continuous visual tokens. To enforce strong semantic grounding and stable latent predictions, the vision head is trained to align the continuous visual embeddings extracted by off-the-shelf visual encoders (e.g., SigLIP-v2tschannen2025siglip2 ). Such alignment can be accomplished by computing the similarities (e.g., L2 distance, cosine similarity) between the visual tokens $V_{o}$ from vision head and visual embeddings $E$ from the pre-trained visial encoders:
@@ 4 Data and Training
JourneyDB-4M sun2023journeydb
LLaVA-v1.5 liu2024llava_1_5
ShareGPT4V chen2024sharegpt4v
Cambrain-7M tong2024cambrian
MMC4 zhu2023mmc4
@@
LaVA-Video-178K zhang2024llava-vid
VSTaR-1M zohar2024vstar
ShareGPT4Video-40K chen2024sharegpt4video
@@
LAION-5B schuhmann2022laion
COYO-700M kakaobrain2022coyo-700m
@@
AnyEdit yu2024anyedit
UltraEdit zhao2024ultraedit
HIVE zhang2024hive
PromptFix yu2024promptfix
@@ 4.3 Training
To mitigate the substantial training costs, we initialize the components of our unified model using readily available pretrained weights. Training either a Multimodal Large Language Model (MLLM) for video understanding or a diffusion model for video generation from scratch necessitates vast computational resources and data. Therefore, the selection of appropriate pretrained models for initialization is a critical step. For the MLLM component, we adopt the weights from VILA lin2024vila . Regarding the diffusion model component, we initialize our model with the weights of the Wan 2.1 text-to-video model wan2025wan . This model was chosen for its compact parameter count of only 1.3B, which still provides a strong foundation for text-to-video generation. We clarify that the original Wan 2.1-1.3B model does not natively support visual editing capabilities, and we enable the diffusion decoder to perform image and video editing tasks via the proposed method mentioned earlier. Moreover, the newly introduced adapter and vision head are initialized using a standard random initialization manner.
@@ 5.2 Experiment Results
We evaluated our model’s long-range video generation performance on the VBench-Long huang2024vbench benchmark, which comprises 16 fine-grained metrics. As shown in Tab. 3, our method achieves state-of-the-art results in Subject Consistency (98.39%), Temporal Flickering (99.87%), Object Class recognition (93.54%), and Temporal Style consistency (25.81%). These results highlight the model’s strong ability to preserve temporal coherence, maintain subject identity, and capture semantic details over extended sequences. Compared with Wan2.1-T2V-1.3B, which reports the highest quality score (85.23%), our model delivers competitive overall performance while surpassing it in 10 out of 16 categories, demonstrating the robustness and generalization ability of our unified video generation framework.
@@
SDv1.5 rombach2022high 0.97 0.38 0.35 0.76 0.04 0.06 0.43
SDv2.1 rombach2022high 0.98 0.51 0.44 0.85 0.07 0.17 0.50
SDXL sdxl 0.98 0.74 0.39 0.85 0.15 0.23 0.55
PixArt-$\alpha$ chen2023pixartalpha 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E 2 dalle2 0.94 0.66 0.49 0.77 0.10 0.19 0.52
DALL-E 3 dalle3 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Chameleon team2024chameleon - - - - - - 0.39
LWM liu2024world 0.93 0.41 0.46 0.79 0.09 0.15 0.47
[…]
Show-o xie2024show 0.95 0.52 0.49 0.82 0.11 0.28 0.53
JanusFlow ma2025janusflow 0.97 0.59 0.45 0.83 0.53 0.42 0.63
@@
Table 3: T2V performance on VBench-Long.
@@
EasyAnimateV5.1 fu2024easyanimate 83.42 85.03 77.01 98.00 97.41 99.19 98.02 57.15 69.48 68.61
MiniMax-Video-01 minimax2024video 83.41 84.85 77.65 97.51 97.05 99.10 99.22 64.91 63.03 67.17
Kling 1.6 kuaishou2025kling 83.40 85.00 76.99 97.40 96.84 98.64 99.13 62.22 64.81 69.70
Wan2.1-T2V-1.3B wan2025wan 83.31 85.23 75.65 97.56 97.93 99.55 98.52 65.19 65.46 67.01
HunyuanVideo kong2024hunyuanvideo 83.24 85.09 75.82 97.37 97.76 99.44 98.99 70.83 60.36 67.56
Gen-3 runway2024gen3 82.32 84.11 75.17 97.10 96.62 98.61 99.23 60.14 63.34 66.82
Vchitect-2.0 (VEnhancer) fan2025vchitect 82.24 83.54 77.06 96.83 96.66 98.57 98.98 63.89 60.41 65.35
CogVideoX1.5-5B yuan2024cogvideox 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02
@@
EasyAnimateV5.1 fu2024easyanimate 89.57 66.85 95.60 77.86 76.11 54.31 23.06 24.61 26.47
MiniMax-Video-01 minimax2024video 87.83 76.04 92.40 90.36 75.50 50.68 20.06 25.63 27.10
Kling 1.6 kuaishou2025kling 93.34 63.99 96.20 81.26 79.08 55.57 20.75 24.51 26.04
Wan2.1-T2V-1.3B wan2025wan 88.81 74.83 94.00 89.20 73.04 41.96 21.81 23.13 25.50
HunyuanVideo kong2024hunyuanvideo 86.10 68.55 94.40 91.60 68.68 53.88 19.80 23.89 26.44
Gen-3 runway2024gen3 87.81 53.64 96.40 80.90 65.09 54.57 24.31 24.71 26.69
Vchitect-2.0 (VEnhancer) fan2025vchitect 86.61 68.84 97.20 87.04 57.55 56.57 23.73 25.01 27.57
CogVideoX1.5-5B yuan2024cogvideox 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30

### v4

@@ 3.1 Overview
Recall that our aim is to build a unified architecture that enables multimodal video understanding and generation simultaneously. To achieve this, we teach multimodal understanding models (e.g.,VILA Lin et al. (2024)) to produce visual tokens as the input of a text-to-video diffusion model with lightweight adapters and a novel three-stage training scheme. In this way, the alignment between the output visual tokens and the diffusion model’s context conditions could be achieved efficiently, without requiring training the whole system from scratch. Fig. 2 presents the overall structure of our Omni-Video. Overall, our model connects an AR-based MLLM model for understanding and a DiT-based diffusion model for visual generation, more details are presented below.
@@ 3.2 Producing Visual Tokens with MLLMs
More specifically, the text token performs standard language modeling, predicting tokens from a vocabulary augmented with four newly introduced special tokens: <BOI>, <EOI>, <BOV>, and <EOV>. Conversely, the vision head maps the MLLMs’ hidden states into a continuous visual tokens. To enforce strong semantic grounding and stable latent predictions, the vision head is trained to align the continuous visual embeddings extracted by off-the-shelf visual encoders (e.g., SigLIP-v2Tschannen et al. (2025)). Such alignment can be accomplished by computing the similarities (e.g., L2 distance, cosine similarity) between the visual tokens $V_{o}$ from vision head and visual embeddings $E$ from the pre-trained visial encoders:
@@ 4 Data and Training
JourneyDB-4M Sun et al. (2023)
LLaVA-v1.5 Liu et al. (2024d)
ShareGPT4V Chen et al. (2024a)
Cambrain-7M Tong et al. (2024a)
MMC4 Zhu et al. (2023b)
@@
LaVA-Video-178K Zhang et al. (2024b)
VSTaR-1M Zohar et al. (2024)
ShareGPT4Video-40K Chen et al. (2024b)
@@
LAION-5B Schuhmann et al. (2022)
COYO-700M Minwoo et al. (2022)
@@
AnyEdit Yu et al. (2024a)
UltraEdit Zhao et al. (2024)
HIVE Zhang et al. (2024a)
PromptFix Yu et al. (2024b)
@@ 4.3 Training
To mitigate the substantial training costs, we initialize the components of our unified model using readily available pretrained weights. Training either a Multimodal Large Language Model (MLLM) for video understanding or a diffusion model for video generation from scratch necessitates vast computational resources and data. Therefore, the selection of appropriate pretrained models for initialization is a critical step. For the MLLM component, we adopt the weights from VILA Lin et al. (2024). Regarding the diffusion model component, we initialize our model with the weights of the Wan 2.1 text-to-video model Wan et al. (2025). This model was chosen for its compact parameter count of only 1.3B, which still provides a strong foundation for text-to-video generation. We clarify that the original Wan 2.1-1.3B model does not natively support visual editing capabilities, and we enable the diffusion decoder to perform image and video editing tasks via the proposed method mentioned earlier. Moreover, the newly introduced adapter and vision head are initialized using a standard random initialization manner.
@@ 5.2 Experiment Results
We evaluated our model’s long-range video generation performance on the VBench-Long Huang et al. (2024) benchmark, which comprises 16 fine-grained metrics. As shown in Tab. 3, our method achieves state-of-the-art results in Subject Consistency (98.39%), Temporal Flickering (99.87%), Object Class recognition (93.54%), and Temporal Style consistency (25.81%). These results highlight the model’s strong ability to preserve temporal coherence, maintain subject identity, and capture semantic details over extended sequences. Compared with Wan2.1-T2V-1.3B, which reports the highest quality score (85.23%), our model delivers competitive overall performance while surpassing it in 10 out of 16 categories, demonstrating the robustness and generalization ability of our unified video generation framework.
@@
SDv1.5 Rombach et al. (2022) 0.97 0.38 0.35 0.76 0.04 0.06 0.43
SDv2.1 Rombach et al. (2022) 0.98 0.51 0.44 0.85 0.07 0.17 0.50
SDXL Podell et al. (2023) 0.98 0.74 0.39 0.85 0.15 0.23 0.55
PixArt-$\alpha$ Chen et al. (2023) 0.98 0.50 0.44 0.80 0.08 0.07 0.48
DALL-E 2 Ramesh et al. (2022) 0.94 0.66 0.49 0.77 0.10 0.19 0.52
DALL-E 3 Betker et al. (2023) 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Chameleon Team (2024) - - - - - - 0.39
LWM Liu et al. (2024c) 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X Ge et al. (2024) 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Show-o Xie et al. (2024) 0.95 0.52 0.49 0.82 0.11 0.28 0.53
JanusFlow Ma et al. (2025) 0.97 0.59 0.45 0.83 0.53 0.42 0.63
@@
Table 3: T2V performance on VBench Huang et al. (2024). Quantitative results are quoted from the official Vbench Leaderboard.
@@
EasyAnimateV5.1 Xu et al. (2025) 83.42 85.03 77.01 98.00 97.41 99.19 98.02 57.15 69.48 68.61
MiniMax-Video-01 MiniMax (2024) 83.41 84.85 77.65 97.51 97.05 99.10 99.22 64.91 63.03 67.17
Kling 1.6 Technology (2025) 83.40 85.00 76.99 97.40 96.84 98.64 99.13 62.22 64.81 69.70
Wan2.1-T2V-1.3B Wan et al. (2025) 83.31 85.23 75.65 97.56 97.93 99.55 98.52 65.19 65.46 67.01
HunyuanVideo Kong et al. (2024) 83.24 85.09 75.82 97.37 97.76 99.44 98.99 70.83 60.36 67.56
Gen-3 25 82.32 84.11 75.17 97.10 96.62 98.61 99.23 60.14 63.34 66.82
Vchitect-2.0 (VEnhancer) Fan et al. (2025b) 82.24 83.54 77.06 96.83 96.66 98.57 98.98 63.89 60.41 65.35
CogVideoX1.5-5B Yang et al. (2024) 82.17 82.78 79.76 96.87 97.35 98.88 98.31 50.93 62.79 65.02
@@
EasyAnimateV5.1 Xu et al. (2025) 89.57 66.85 95.60 77.86 76.11 54.31 23.06 24.61 26.47
MiniMax-Video-01 MiniMax (2024) 87.83 76.04 92.40 90.36 75.50 50.68 20.06 25.63 27.10
Kling 1.6 Technology (2025) 93.34 63.99 96.20 81.26 79.08 55.57 20.75 24.51 26.04
Wan2.1-T2V-1.3B Wan et al. (2025) 88.81 74.83 94.00 89.20 73.04 41.96 21.81 23.13 25.50
HunyuanVideo Kong et al. (2024) 86.10 68.55 94.40 91.60 68.68 53.88 19.80 23.89 26.44
Gen-3 25 87.81 53.64 96.40 80.90 65.09 54.57 24.31 24.71 26.69
Vchitect-2.0 (VEnhancer) Fan et al. (2025b) 86.61 68.84 97.20 87.04 57.55 56.57 23.73 25.01 27.57
CogVideoX1.5-5B Yang et al. (2024) 87.47 69.65 97.20 87.55 80.25 52.91 24.89 25.19 27.30